OpenStudy (anonymous):

marthas garden is 5 m wide and 8 m long. next year she plans to double the gardens area by increasing the width and lengthe by the same number of meters. what will be the new dimensions of the garden?

jimthompson5910 (jim_thompson5910):

Original Area: A = LW A = 5*8 A = 40 New Area A = LW A = (5+x)*(8+x) 80 = 40 + 13x + x^2 x^2 + 13x + 40 - 80 = 0 x^2 + 13x - 20 = 0

jimthompson5910 (jim_thompson5910):

Solve for x to find the amount added to each dimension

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks for the help i got a 90% on my quiz